OpenStudy (anonymous):

Let sin p=.71 What is cos(2p)

OpenStudy (anonymous):

is it 0.71 ??

OpenStudy (anonymous):

in the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\cos2\pi= 1- 2\sin^2 \pi = 1-2(0.71)^2=1-2\times0.5041=1-1.0082=-0.0082\]

OpenStudy (anonymous):

thank you so much ;D